OpenStudy (anonymous):

A tugboat goes upstream in 15 hours. The rturn trip downstream takes 10 hours. Find the speed of th tugboat without a current and the speed of current

OpenStudy (anonymous):

i have 120/15=8 120/10=12

OpenStudy (anonymous):

how long is the trip?

OpenStudy (anonymous):

Like in miles or kilometers or something

OpenStudy (anonymous):

it only ask for tugboat is ? mph and speed of the current is ? mph

OpenStudy (anonymous):

Yeah, you need the distance, is that all the info it gave you?

OpenStudy (anonymous):

120/15=8 120/10=12 upstram=8 downstream=12

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

2b=20mph b=10mph b+c=12 10+c=12 10-12 c=2

OpenStudy (anonymous):

i hope i got it correct

OpenStudy (anonymous):

I'm not sure what to do if it didnt give you the distance of the trip. It doesnt look like enough info to do anything

OpenStudy (anonymous):

ikw so thats what i cam up with

OpenStudy (anonymous):

You're sure thats all it gave you? Where did you get 120 from?

OpenStudy (anonymous):

ok im apologize iit has a tugboat goes 120 upstream

OpenStudy (anonymous):

so the distance of the trip is 120 miles?

OpenStudy (anonymous):

in 15hours.

OpenStudy (anonymous):

yes i had to reread...

OpenStudy (anonymous):

Alright, well distance=speed*time so let x be the speed of the tugboat without current and let y be the speed of the current. You're gonna have 2 equations. 120=(x+y)10 120=(x-y)15 Now you distribute the numbers outside the parentheses and then solve it as a system of equations

OpenStudy (anonymous):

so my answer is wrong?

OpenStudy (anonymous):

what did you get for the speed of the current and for the speed of the boat?

OpenStudy (anonymous):

10 and 2

OpenStudy (anonymous):

2mph=speed of current 10mph=speed of boat without current So you're right

OpenStudy (anonymous):

oh ok cool thanks for chcking it